OpenStudy (anonymous):

if u = -1-sqrt 3i , write vector of u in polar form

OpenStudy (anonymous):

you need two things, \(r\) and \(\theta\)

OpenStudy (anonymous):

\(r\) is easiest, it is \(r=\sqrt{a^2+b^2}\) which in your case it \[r=\sqrt{1^2+\sqrt{3}^2}=\sqrt{1+3}=\sqrt4=2\]

OpenStudy (anonymous):

then you need \(\theta\) where \[\cos(\theta)=\frac{a}{r}=-\frac{1}{2}\] and \[\sin(\theta)=\frac{b}{r}=-\frac{\sqrt3}{2}\]

OpenStudy (anonymous):

should be clear what \(\theta\) is, then you are done once you write it as \[r\left(\cos(\theta)+i\sin(\theta)\right)\]

OpenStudy (anonymous):

hi thanks for your help, so i'm given an argand diagram should i plot on it the -srt3/2 and -1/2 ?

OpenStudy (anonymous):

i guess so is an "argand diagram" the same as the complex plane?

OpenStudy (anonymous):

oh, yes it is ok right

OpenStudy (anonymous):

yeah so are the point above that i mentioned should be drawn?